OpenStudy (anonymous):

1. Suppose a and b are real numbers where a > 0 and b < 0. Please give me the answer or a really well explanation a. Is –a + b positive or negative. Explain how you know. b. Is a – b positive or negative? Explain how you know. c. Is b – a positive or negative. Explain how you know

OpenStudy (mathstudent55):

a. Since a is positive, -a is negative. b is negative. The sum of two negative numbers is negative, so -a + b is negative. b. a - b = a + (-b). a is positive, and b is negative, so -b is positive. The sum of two positive numbers is positive. c. b - a = b + (-a). b is negative. a is positive, so -a is negative. The sum of two negative numbers is negative.

OpenStudy (anonymous):

thanks a bunch c:

OpenStudy (mathstudent55):

wlcm